You want to know whether national disasters significantly affect anxiety. You use an established anxiety scale...

  1. You want to know whether national disasters significantly affect anxiety. You use an established anxiety scale with μ = 50, σ = 45). You sample N = 87 individuals from San Francisco right after a major earthquake, and the sample has M = 70. Identify the independent (+levels) and dependent variables. Does the independent variable significantly affect the dependent variable? What do you conclude? Use ɑ = 0.05.
Answer #1

Sol:

Ho:μ = 50

Ha:μ not = 50

alpja=0.05

Independent variable-national disasters

dependent variable-anxiety

z=xbar-mu/s/sqrt(n)

z=(70-50)/(45/sqrt(87)

z=4.145502

p value in excel is

=NORM.S.DIST(-4.145502,TRUE)

=1.69535E-05

=2*1.69535E-05
P=3.3907E-05

p=0.000034

p<0.05

Reject Ho

Accept Ha

Conclusion:
There is suffcient statistical evidence at 5% level of signiifcance to conlcude that national disasters significantly affect anxiety